5.) A car traveling at 15 m/s passes a van that is starting from rest with an acceleration of 3 m/s². (a) How long will it take the van to catch up with the car? (b) How far has the car traveled when the van catches up with it?
1
Expert's answer
2019-08-26T10:41:37-0400

The equation of the path traveled by the car


"S_c=u*t;"

u=15 m/s;

The equation of the path traveled by the van


"S_v=\frac{a*t^2}{2};"

a= 3 m/s^2;

a) When the truck catches up with the car:


"S_c=S_v;
\\u*t=\frac{a*t^2}{2};
\\ t=\frac{2*u}{a}
\\t=10s;"

b)substitute time into the equation Sc:


"S_c=u*t;
\\S_c=150 m"